The Smelting Department of Kiner Company has the following production and cost data for November.

Production: Beginning work in process 2,800 units that are 100% complete as to materials and 29% complete as to conversion costs; units transferred out 9,700 units; and ending work in process 8,200 units that are 100% complete as to materials and 44% complete as to conversion costs.

Compute the equivalent units of production for (a) materials and (b) conversion costs for the month of November.

   

Materials

 

          Conversion Costs

Total equivalent units  
Answer #1
Concepts and reason

The concept used to solve the problem is Process Costing.

Process Costing: Under the method of process costing, the cost is assigned to production units in companies producing huge quantities of similar products. This operation is used for determining product cost at each stage of the process or stage of the manufacturing.

Fundamentals

Work-in-progress (WIP) Inventory: The inventory at which some percentage of production activity has been performed is WIP inventory. It is the inventory which is not 100% complete for final consumption.

Raw material inventory: The inventory which is purchased from the supplier for the purpose of using it as an ingredient while manufacturing the product. The inventory which is directly issued for production is direct material and which is indirectly issued to the production process are indirect material. The inventory at which no production activity is performed is raw material inventory. It is also known as stores ledger control account because it is kept in the storeroom and will be issued for production from stores.

Conversion Cost: This is the cost which is incurred during the process of transferring the raw material inventory into the finished product. It includes direct labor and overheads that is used to convert the raw material into the finished goods.

Direct labor: Workers or employees directly involved in the manufacture of goods and services are called the direct labor. Direct labor can be directly traceable to a specific product or cost center. In other words, the cost of the workers which is directly involved in the transformation of the raw material into the finished product.

Manufacturing Overhead: These are the expenditures which are directly associated with the manufacturing of goods or services other than the direct labor and direct material. All indirect material, indirect labor, indirect overheads come under this category.

Equivalent units: It is the amount of work that has been done by the producer. These are the partly completed units till the end of the accounting period which has been presented in terms of completely produced units.

(a)

Compute the equivalent unit produced in terms of material:

Particulars
Units transferred out
Ending Inventory
Equivalent units of production
Units
9,700
8,200
17,900

Working notes:

Compute the ending inventory:

EndingInventory=EndingInventory×percentofcompletionintermsofmaterial=8,200units×100%=8,200units\begin{array}{c}\\{\rm{Ending}}\,{\rm{Inventory}} = {\rm{Ending}}\,{\rm{Inventory}} \times {\rm{percent}}\,{\rm{of}}\,{\rm{completion}}\,{\rm{in}}\,{\rm{terms}}\,{\rm{of}}\,{\rm{material}}\\\\ = 8,200\,{\rm{units}} \times 100\% \\\\ = 8,200\,{\rm{units}}\\\end{array}

Compute the equivalent units of production:

Equivalentunitsofproduction=Unitstransferredout+EndingInventory=9,700units+8,200units=17,900units\begin{array}{c}\\{\rm{Equivalent}}\,{\rm{units}}\,{\rm{of}}\,{\rm{production}} = {\rm{Units}}\,{\rm{transferred}}\,{\rm{out}} + {\rm{Ending}}\,{\rm{Inventory}}\\\\ = 9,700\,{\rm{units}} + 8,200\,{\rm{units}}\\\\ = 17,900\,{\rm{units}}\\\end{array}

(b)

Compute the equivalent unit produced in terms of conversion cost:

Particulars
Units transferred out
Ending Inventory
Equivalent units of production
| Units
9.700
3,608
13,308

Working notes:

Compute the ending inventory:

EndingInventory=EndingInventory×percentofcompletionintermsofConversioncost=8,200units×44%=3,608units\begin{array}{c}\\{\rm{Ending}}\,{\rm{Inventory}} = {\rm{Ending}}\,{\rm{Inventory}} \times {\rm{percent}}\,{\rm{of}}\,{\rm{completion}}\,{\rm{in}}\,{\rm{terms}}\,{\rm{of}}\\\\\,{\rm{Conversion}}\,{\rm{cost}}\\\\ = 8,200\,{\rm{units}} \times 44\% \\\\ = 3,608\,{\rm{units}}\\\end{array}

Compute the equivalent units of production:

Equivalentunitsofproduction=Unitstransferredout+EndingInventory=9,700units+3,608units=13,308units\begin{array}{c}\\{\rm{Equivalent}}\,{\rm{units}}\,{\rm{of}}\,{\rm{production}} = {\rm{Units}}\,{\rm{transferred}}\,{\rm{out}} + {\rm{Ending}}\,{\rm{Inventory}}\\\\ = 9,700\,{\rm{units}} + 3,608\,{\rm{units}}\\\\ = 13,308\,{\rm{units}}\\\end{array}

Ans: Part a

Equivalent units produced in term of material 17,900 units.

Part b

Equivalent units of production in terms of conversion cost 13,308 units.